OpenStudy (anonymous):

Is x^2-4x+24 factorable?

hero (hero):

Any quadratic is factorable

hero (hero):

Even if the book says it isn't

OpenStudy (anonymous):

Ok thanks what does it factor to then?

OpenStudy (curry):

hero go on chat

hero (hero):

That would probably take a bit of work to factor but you have to find two numbers a*c and then those same two numbers would add up to b = -4

hero (hero):

To help, a*c = 1*24 = 24

OpenStudy (anonymous):

a is given as 1, what would i do knowing that?

hero (hero):

a = (b/2+d) c = (b/2-d)

hero (hero):

knowing all of that....we can do this: b = a+c = (b/2+d)+(b/2-d)

hero (hero):

In other words....

hero (hero):

We can multiply that out to find d

OpenStudy (anonymous):

Ok thank you so much!

hero (hero):

I'm not done yet

hero (hero):

basically, we know that d will equal + or - 4

hero (hero):

so....

hero (hero):

I just lost my train of thought

hero (hero):

24 = (-4/2+d)(-4/2-d)

hero (hero):

That's the right equation

hero (hero):

Well, the solution is complex, so I guess they consider than unfactorable